OpenStudy (anonymous):

What is the equation of the graph that represents the parent function f(x4) stretched vertically by a factor of 2, and then shifted up 3 spaces.

OpenStudy (anonymous):

A. g(x) = 2x^4 - 3 B. g(x) = 2x^4 + 3 C. g(x) = 2(x + 3)^4 D. g(x) = 2(x^4 + 3)

OpenStudy (anonymous):

the answer is B:)